How do you find the slope of a line going through the given pair of points.
(1/8,1/9) and (3/8,2/9)

Answer by checkley77(12844)

(1/8,1/9) and (3/8,2/9)
Slope=(y2-y1)/(x2-x1)
Slope=(2/9-1/9)/(3/8-1/8)
Slope=(1/9)/(2/8)
Slope=(1/9)*(8/2) (invert the denominator & multiply)
Slope=8/18
Slope=4/9 ans.